OpenStudy (anonymous):

By using limit definition: prove x-->2 lim x^2 = 4 How to get epsilon for |x-2|.|x+2| < epsilon. why would i put |x-2| < 1!!

OpenStudy (anonymous):

\[ |x-2|\lt \delta \implies |x^2-4|<\epsilon\\ |x-2|\lt \delta \implies |x-2||x+2|<\epsilon\\ |x-2|\lt \delta \implies |x-2|<\frac{\epsilon}{|x+2|}\\ \]Ideally you would want to say \[ \delta = f(\epsilon) = \frac{\epsilon}{|x+2|} \]However, you don't get to know what \(x\) will be, so you have to be a bit more clever.

OpenStudy (anonymous):

What you can do in this case though, is to find some minimum delta range that you can always fallback on. In this case, we will let \(\delta=1\) be the fallback.

OpenStudy (anonymous):

yea why 1

OpenStudy (anonymous):

So we really only have to worry about cases where \(2-1\leq x \leq 2+1\). This means that \(1\leq x \leq 3\).

OpenStudy (anonymous):

To be honest, it could be any number grater than \(0\).

OpenStudy (anonymous):

We could have chosen \(1/2000\). I chose \(1\) because I think it will be an easy number

OpenStudy (anonymous):

so i can pick 5 or 10 my text always choose it to be 1

OpenStudy (anonymous):

Yes, you can pick 5 or 10 if you want

OpenStudy (anonymous):

Typically you want to pick a sufficiently small number.

OpenStudy (anonymous):

Since \(1\) is easy to work with, a lot of times \(1\) will be picked.

OpenStudy (anonymous):

However, do you mind if I actually just step back a bit and explain better what I am trying to accomplish?

OpenStudy (anonymous):

nope ,but i know the rest i just wondered that text just picked one!!

OpenStudy (anonymous):

Do you understand why they had to pick a number?

OpenStudy (anonymous):

nope :D

OpenStudy (anonymous):

Okay so let's go back to this for a moment:\[ \delta = f(\epsilon) = \frac{\epsilon}{|x+2|} \]We don't really know what \(x\) is. We want to be safe. The safest we can possibly is to have \(\delta\) as small as possible. Thus what we want to say is: \[ \delta = f(\epsilon) = \min\left(\frac{\epsilon}{|x+2|}\right) \]

OpenStudy (anonymous):

It is always safer to make \(\delta\) smaller. The smaller it is, the smaller our \(\epsilon\) range will be.

OpenStudy (anonymous):

Does this make sense?\[ \min\left(\frac{\epsilon}{|x+2|}\right) = \frac{\epsilon}{\max(|x+2|)} \]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

The minimum is where the denominator is maximum. The problem now is that there is no \(\max\) to \(|x+2|\). It could go to \(\infty\).

OpenStudy (anonymous):

We can't pick \(\epsilon/\infty\) because \(\epsilon/\infty = 0 \) and we must pick \(\delta\) to be greater than \(0\).

OpenStudy (anonymous):

So we have to pick a cutoff point.

OpenStudy (anonymous):

In the book, they said \(\delta \le 1\). Since we get to chose what \(\delta\) is, we can always create a restriction like this.

OpenStudy (anonymous):

The result of this is that \(|x-2|<1\implies 1\). This means that \(\max(|x+2|) = 5\), and so: \[ \delta = \min\left(1,\frac{\epsilon }{5}\right) \] Would you like me to show what happens when we say \(\delta \le 5\)?

OpenStudy (anonymous):

Does this is related with that the maximum is when x+2=1 ,so x =-1 ??

OpenStudy (anonymous):

\[ |x-2|<5 \implies 2-5<x<2+5\implies -3<x<7\implies 0<|x+2|<9 \]

OpenStudy (anonymous):

\[ \implies |x+2|<9 \]

OpenStudy (anonymous):

-1 and 9

OpenStudy (anonymous):

Well, the absolute values ensures it is greater that \(0\).

OpenStudy (anonymous):

Or perhaps I should say greater than or equal

OpenStudy (anonymous):

OK ,but does the one is related to the maximum is when x =-1

OpenStudy (anonymous):

That max of what?

OpenStudy (anonymous):

epsilon / x+2

OpenStudy (anonymous):

It will max out when you approach \(-2\)

OpenStudy (anonymous):

oh uh... I got wrong. So that is write for prove to choose 1 ,but in real i would choose large number wright?

OpenStudy (anonymous):

You want to chose \(\delta\) to be small, actually.

OpenStudy (anonymous):

Unfortunately you can't chose \(\delta=0\).

OpenStudy (anonymous):

Certainly, is would be safer to choose \(10^{-10000000}\) than choosing \(1=10^0\).

OpenStudy (anonymous):

However, if you chose that, then you'd end up with: \[ |x-2|<10^{-1000000} \implies 2-10^{-1000000}<x<2+10^{-1000000}\\ \implies |x+2| < 4+10^{-1000000} \]See... It just makes it more complicated. It is easier to work with \(1\).

OpenStudy (anonymous):

And since \(1\) is good enough, then might as well chose \(1\). It really is that simple.

OpenStudy (anonymous):

x-2 < min (e/x+2) min when x ---> infinity and the left side grows also ????

OpenStudy (anonymous):

what is wrong with that i am confused

OpenStudy (anonymous):

Remember what epsilon really is....

OpenStudy (anonymous):

\[ |f(x)-L|<e \]

OpenStudy (anonymous):

So you are saying: \[ |x-2| \leq \min\left(\frac{|x^2-4|}{|x+2|}\right)\lt \min\left(\frac{\epsilon}{|x+2|}\right) \]

OpenStudy (anonymous):

So yes, the left site will grow as well, because \(\epsilon\) range will have to grow.

OpenStudy (anonymous):

As \(x\) departs from \(2\), \(f(x)\) departs from \(4\).

OpenStudy (anonymous):

Well, potentially, anyway

OpenStudy (anonymous):

OK . I think I got it now. Thank you very much man .